# Refined asymptotics for the Cauchy problem for the fast $p$-Laplace evolution equation

### v2

@@ 1.2 The very fast diffusion range: $1<p\leq p_{c}$
The long-time behaviour of solutions to (CPLE) in the very fast diffusion range $1<p\leq p_{c}$ remains poorly understood. In fact, this regime presents several interesting open questions, many of which are still unanswered after more than 50 years of deep study. While solutions exist when $0\leq{{u_{0}}}\in\mathrm{L}^{1}_{\textrm{loc}}({\mathbb{R}^{N}})$ for every $p>1$, see [42], several key properties of them that hold in the good range, are lost in the very fast diffusion regime, see Section 3. Let us begin with the failure of the mass conservation. As already mentioned, the value $p=p_{c}$ is critical for the conservation of mass, i.e. (1.1), of solutions that are in $\mathrm{L}^{1}({\mathbb{R}^{N}})$. It is known that in this case solutions whose initial datum is in $\mathrm{L}^{1}({\mathbb{R}^{N}})$ still conserve mass, see for instance the recent survey [74]. However, when $1<p<p_{c}$ the mass is not conserved anymore. More surprisingly, for a large class of data (i.e., ${{u_{0}}}\in L^{r}({\mathbb{R}^{N}})$ with $r=n(2-p)/p$ and/or ${{u_{0}}}(x)\leq c{V_{D}}(x)$ for some $D>0$) the corresponding solution extinguishes at $T>0$ (i.e., $u(t,\cdot)=0$, for all $t>T$ a.e. in ${\mathbb{R}^{N}}$), cf. [72, Chapter 11].
@@
For $p=p_{Y}$, in [50] the authors find an explicit solution obtained by the separation of variables (see [50, (8.11)]) that can be written as
@@
In this case, by using the results of [50] and the more geometric setting of [33], it can be proven that $U$ is an attractor for a certain class of solutions when $p=p_{Y}$. However, a broad picture is far more complex: due to [50, Theorem 8.2] it is known that for any $p\in\left(1,p_{c}\right)$ a family of self-similar solutions with different tails from $U$ and $\mathcal{B}_{D,T}$ exist. To the best of our knowledge the following natural question has not been addressed before:
What is the basin of attraction of the pseudo-Barenblatt solution in the very fast diffusion regime? $\begin{array}[]{c}\textit{What is the basin of attraction of the pseudo-Barenblatt solution in the very fast diffusion regime? }\end{array}$ (Q-3)
@@
Note that condition (1.29) is satisfied in low dimensions ($61<N<$ 6) for all $p\in(1,2)$, while in high dimensions ($N\geq 6$) for $p$ close to $1$ and close to $2$, precisely for $p\in(1,p_{1})\cup(p_{2},2)$ where
@@
Range(s) of validity and open questions. In the case $62\leq N<$ 6, the entire range $p\in(1,2)\setminus\{\frac{3}{2}\}$ is covered. When $N\geq 6$, condition (1.29) is equivalent to $p\in(1,p_{1})\cup(p_{2},2)$ with $p_{1}$, $p_{2}$ as in (1.30). Therefore, Theorem 4 not only extends the range of admissible values of $p$ close to $2$ from $(p_{c},2)$ to $(p_{2},2)$, but also covers a very singular range close to $1$. The range $[p_{1},p_{2}]$ is not covered in high dimensions and we believe that our methods cannot be extended to this range. This is a difficult and intriguing open question that requires new ideas.
@@
Convergence of radial derivatives in the very fast diffusion range. As in the good fast diffusion range, when (Q-3) is (at least partially) resolved, it is natural to ask whether radial derivatives of solutions converge towards those of the pseudo-Barenblatt profile, i.e. question (Q-2) in the context of the very fast diffusion range. Generically, once a result as Theorem 5 is proven, this question could be settled by applying regularity results and interpolation inequalities (as in [10]). However, as it is explained in Remark 1.4, in the present setting good regularity results are missing. We have, therefore, decided to use a different technique which uses a correspondence between radial derivatives of (R-CPLE) (or equivalently (CPLE)) with those of a weighted diffusion equation of porous medium type, see Section 6. In our analysis there appear again the exponent $p_{Y}$, defined in (1.26), this time as a threshold exponent. We have decided to call this exponent after Yamabe since, in this case, the radial derivative of (CPLE) solves the equation $\partial_{t}\Phi=\Delta\Phi^{\frac{N-2}{N+2}}$ (see Section 6 for more information). This equation is related to the Yamabe flow for a conformally flat metric, see [33, 30, 80, 10]. The value $p_{Y}$ is also a sharp threshold for the gradient regularity of solutions when no extra assumptions are imposed, see Section 3.3 and [41, Section 21.3]. We stress also that $\max\{p_{Y},p_{2}\}<p_{c}$, but the relation between $p_{Y}$ and $p_{2}$ depends on the dimension. Recall that $p^{\prime}=p/(p-1)$. Let us now present our main result.
@@
We notice again that exponent $p_{Y}$ plays an important role in any dimension, while $p_{1}$ and $p_{2}$ appear only when $N\geq 6$. Therefore it makes sense in our analysis to distinguish between these cases. When $62\leq N<$ 6, we recall that Theorem 4 gives the sufficient condition for an answer to question (Q-3), and under the additional assumption (1.33), rates are provided in Theorem 5. Such conditions are verified in the radial case, see Section 9. When it comes to the to convergence of radial derivatives in relative error, i.e. Theorem 6, the results hold up to $p_{Y}$. The second important exponent is $p=\frac{3}{2}$, when the entropy (at least in the form (1.12)) is not defined anymore and we cannot adapt the results of Theorem 5. It is important to notice that $p_{c}$ could be smaller than $\frac{3}{2}$ (in $N=2$), equal to $\frac{3}{2}$ (in $N=3$), or larger than $\frac{3}{2}$ (in $N\in\{4,5\}$), which complicates the graphical representation. We also stress that, for any $D_{2},D_{1}>0$ we have $V_{D_{2}}-V_{D_{1}}\in\mathrm{L}^{1}(\mathbb{R}^{N})$ for any $21<p<$ 2 and $N\in\{2,3,4,5\}$. In such a case, and under assumption (1.31) the relative entropy ${\mathcal{E}}[v|{V_{D}}]\in\mathrm{L}^{\infty}(\tau_{0},\infty)$ for any $\tau_{0}>0$. The conservation of relative mass, i.e. property (ii) in Theorem 5, is expected to hold under other assumptions of Theorem 5, as it holds for solutions whose initial data is radial-decreasing, see Section 9. It is unknown whether the Fisher information is bounded along the flow, or not. However, this property holds in the case of radially decreasing initial data, see Section 9. Inequality (1.15), i.e. the fact that the Fisher information controls the entropy, can be obtained by comparison with the Hardy–Poincaré inequality, see Proposition 5.9, under the additional assumption (1.33). When $N\geq 6$, the situation is similar but a major difference: the exponents $p_{1}$ and $p_{2}$ appears and $V_{D_{2}}-V_{D_{1}}\in\mathrm{L}^{1}(\mathbb{R}^{N})$ holds only when $p\in\left(1,p_{1}\right)\cup\left(p_{2},2\right)$. Consequently, the conservation of relative mass and the boundedness of the entropy, could hold only in this last range. At the same time, when $p\in\left[p_{1},p_{2}\right]$ the Hardy–Poincaré inequality of Proposition 5.9 does not hold any more. Instead, when $p\in\left(p_{1},p_{2}\right)$ a Hardy-type inequality holds, see [28]. In Table 2 we resume such properties for $N\geq 7$. We notice that the acronym “RDDI” stand for “radial, decreasing whose derivatives satisfy a specific inequality”. Such a class of data will be fully described in Section 9. The case $N=6$ is very similar with the only caveat that $p_{2}=p_{Y}=\frac{3}{2}$.
@@
Range $V_{D_{2}}-V_{D_{1}}\in\mathrm{L}^{1}(\mathbb{R}^{N})$ ${\mathcal{E}}[v(\tau)|{V_{D}}]$ ${\mathcal{I}}[v(v)|{V_{D}}]$ $\tfrac{\,{\rm d}}{\,{\rm d}\tau}{\mathcal{E}}[v(\tau)]=-{\mathcal{I}}[v(v)]$ Inequality (1.15) ${\mathcal{I}}[v(v)|{V_{D}}]\geq c\,{\mathcal{E}}[v(\tau)|{V_{D}}]$ Convergence in relative error for radial derivatives
@@ 2 Results in the original variables
$\ \Big{\|}\frac{\partial_{r}u(t,\cdot)}{\partial_{r}\mathcal{B}_{M}(t,\cdot)}-1\Big{\|}_{L^{\infty}(\mathbb{R}^{N})}\leq k_{\star}\,t^{-\lambda}\,\quad\forall\,t\geq t_{\star}\,,$
@@ 3.2 Existence and uniqueness
We say that $u$ is a non-negative weak solution to problem (CPLE) on $(0,\infty)\times\mathbb{R}^{N}$ for $21<p<$ 2 with non-negative initial data ${{u_{0}}}\in\mathrm{L}^{1}_{{\rm loc}}(\mathbb{R}^{N})$ if $u\in\mathrm{L}^{p}((0,\infty);W^{1,p}_{\rm loc}(\mathbb{R}^{N}))\cap C((0,\infty);\mathrm{L}^{1}_{{\rm loc}}(\mathbb{R}^{N}))$ and
@@ 3.3 Boundedness and regularity of solutions
Here we shall mainly focus on the concept of local weak solution (not necessarily non-negative ones) which differ from our definition above by assumptions on the integrability properties of $u$, namely for a solution defined on $(0,T)\times\Omega$ one asks typically that $u\in\mathrm{L}^{\infty}_{\rm loc}((0,T);\mathrm{L}^{2}_{\rm loc}(\Omega))$ and $|\nabla u|\in\mathrm{L}^{p}_{\rm loc}((0,T);\mathrm{L}^{p}_{\rm loc}(\Omega))$. To the best of our knowledge, these kind of solutions have been studied first in [37] where continuity of $\nabla u$ has been proven, with an explicit modulus of continuity, for $p>\max\{1,p_{Y}\}=\max\{1,\frac{2N}{N+2}\}$. In the same range of $p$, the Hölder continuity of the gradient has been obtained in [39] (with some mistakes in the computations, as it was pointed out in [79], which have been fully solved in [38]), see also the review [76]. The threshold value $p_{Y}=\frac{2N}{N+1}$ is sharp for the gradient regularity under no extra assumptions on the solution. Indeed, when $p\leq p_{Y}$ weak solutions are not bounded, see the discussion [41, Section 21.3]. However, in the whole range $21<p<$ 2 it is known, that bounded weak solutions are Hölder continuous, see [78, 77]. In our case, it is known that solutions to (CPLE) are bounded provided the initial datum ${{u_{0}}}\in\mathrm{L}^{q}_{\rm loc}(\mathbb{R}^{N})$ for $q>N\frac{(2-p)}{p}$, see for instance [42, Section III] and [21, Theorem 2.1]. In conclusion, when $p\in(p_{Y},2)$ and the initial datum is integrable enough, weak solutions to (CPLE) are bounded and, therefore, the function $(t,x)\mapsto u(t,x)$ is $C^{1,\alpha}_{\rm loc}((0,\infty)\times\mathbb{R}^{N})$. However, when $p<p_{c}$, the coefficient $\alpha$ may depend on the function itself other than on $p$ and $N$.
In the present paper, we shall also consider the case $p\leq p_{Y}$, so let us comment on how to obtain the $C^{1,\alpha}_{\rm loc}(\mathbb{R}^{N})$ regularity for solutions in this case. Let us stress, however, that the following reasoning holds in the whole range $21<p<$ 2. The main idea is to use the concept of viscosity solutions (see [53] for a precise definition). In our case, when a weak solution is continuous, then it is also a viscosity solution, see for instance [53, 70, 45, 44]. Since we shall consider only bounded weak solutions (which are Hölder continuous), the above discussion proves that solutions to problem (CPLE) are, indeed, viscosity solutions. The main advantage of employing this notion of solutions is that, in the last decade, there has been a growing interest in obtaining regularity results for viscosity solutions to equations related to (CPLE), see for instance [45] for a detailed bibliography. For what concerns our investigation, the needed result is [51, Theorem 1.1], where the authors prove the function $(t,x)\mapsto u(t,x)$ is $C^{1,\alpha}_{\rm loc}((0,\infty)\times\mathbb{R}^{N})$ for some $\alpha>0$, which depends on $p$, $N$ and the solution itself.
@@
Since the uniform convergence in relative error is related to Harnack inequalities, let us conclude this subsection with some considerations on them. The problem of obtaining a precise form of those inequalities has been a long-standing quest. Indeed, in this nonlinear setting, the intrinsic cylinders depend on the solution itself, showing several differences between the case $p_{c}<p<2$ and $1<p\leq p_{c}$. In the good range, we refer to the paper [40], while in the whole range $21<p<$ 2, it has been proven in [21]. The Harnack inequalities considered in those two papers are valid for local solutions, i.e., no assumption on boundary data is made. For boundary Harnack inequalities, we refer to [61]. Nowadays, several related results are available, see the monograph [40] and references therein.

### v3

@@ 1.2 The very fast diffusion range: $1<p\leq p_{c}$
The long-time behaviour of solutions to (CPLE) in the very fast diffusion range $1<p\leq p_{c}$ remains poorly understood. In fact, this regime presents several interesting open questions, many of which are still unanswered after more than 50 years of deep study. While solutions exist when $0\leq{{u_{0}}}\in\mathrm{L}^{1}_{\textrm{loc}}({\mathbb{R}^{N}})$ for every $p>1$, see [43], several key properties of them that hold in the good range, are lost in the very fast diffusion regime, see Section 3. Let us begin with the failure of the mass conservation. As already mentioned, the value $p=p_{c}$ is critical for the conservation of mass, i.e. (1.1), of solutions that are in $\mathrm{L}^{1}({\mathbb{R}^{N}})$. It is known that in this case solutions whose initial datum is in $\mathrm{L}^{1}({\mathbb{R}^{N}})$ still conserve mass, see for instance the recent survey [75]. However, when $1<p<p_{c}$ the mass is not conserved anymore. More surprisingly, for a large class of data (i.e., ${{u_{0}}}\in L^{r}({\mathbb{R}^{N}})$ with r=n(2p)/pr=n(2-p)/p and/or ${{u_{0}}}(x)\leq c{V_{D}}(x)$ for some $D>0$) the corresponding solution extinguishes at $T>0$ (i.e., $u(t,\cdot)=0$, for all $t>T$ a.e. in $\mathbb{R}^{N}$}), cf. [73, Chapter 11].
@@
For $p=p_{Y}$, in [51] the authors find an explicit solution obtained by the separation of variables (see [51, (8.11)]) that can be written as
@@
In this case, by using the results of [51] and the more geometric setting of [34], it can be proven that $U$ is an attractor for a certain class of solutions when $p=p_{Y}$. However, a broad picture is far more complex: due to [51, Theorem 8.2] it is known that for any $p\in\left(1,p_{c}\right)$ a family of self-similar solutions with different tails from $U$ and $\mathcal{B}_{D,T}$ exist. To the best of our knowledge the following natural question has not been addressed before:
[…]
@@
Note that condition (1.29) is satisfied in low dimensions ($1<N<6$) for all $p\in(1,2)$, while in high dimensions ($N\geq 6$) for $p$ close to $1$ and close to $2$, precisely for $p\in(1,p_{1})\cup(p_{2},2)$ where
@@
Range(s) of validity and open questions. In the case $2\leq N<6$, the entire range $p\in(1,2)\setminus\{\frac{3}{2}\}$ is covered. When $N\geq 6$, condition (1.29) is equivalent to $p\in(1,p_{1})\cup(p_{2},2)$ with $p_{1}$, $p_{2}$ as in (1.30). Therefore, Theorem 4 not only extends the range of admissible values of $p$ close to $2$ from $(p_{c},2)$ to $(p_{2},2)$, but also covers a very singular range close to 11. The range $p_{1},p_{2}]$ is not covered in high dimensions and we believe that our methods cannot be extended to this range. This is a difficult and intriguing open question that requires new ideas.
@@
Convergence of radial derivatives in the very fast diffusion range. As in the good fast diffusion range, when (Q-3) is (at least partially) resolved, it is natural to ask whether radial derivatives of solutions converge towards those of the pseudo-Barenblatt profile, i.e. question (Q-2) in the context of the very fast diffusion range. Generically, once a result as Theorem 5 is proven, this question could be settled by applying regularity results and interpolation inequalities (as in [10]). However, as it is explained in Remark 1.4, in the present setting good regularity results are missing. We have, therefore, decided to use a different technique which uses a correspondence between radial derivatives of (R-CPLE) (or equivalently (CPLE)) with those of a weighted diffusion equation of porous medium type, see Section 6. In our analysis there appear again the exponent $p_{Y}$, defined in (1.26), this time as a threshold exponent. We have decided to call this exponent after Yamabe since, in this case, the radial derivative of (CPLE) solves the equation $\partial_{t}\Phi=\Delta\Phi^{\frac{N-2}{N+2}}$ (see Section 6 for more information). This equation is related to the Yamabe flow for a conformally flat metric, see [34, 30, 81, 10]. The value $p_{Y}$ is also a sharp threshold for the gradient regularity of solutions when no extra assumptions are imposed, see Section 3.3 and [42, Section 21.3]. We stress also that $\max\{p_{Y},p_{2}\}<p_{c}$, but the relation between $p_{Y}$ and $p_{2}$ depends on the dimension. Recall that $p^{\prime}=p/(p-1)$. Let us now present our main result.
@@
We notice again that exponent $p_{Y}$ plays an important role in any dimension, while $p_{1}$ and $p_{2}$ appear only when $N\geq 6$. Therefore it makes sense in our analysis to distinguish between these cases. When $2\leq N<6$, we recall that Theorem 4 gives the sufficient condition for an answer to question (Q-3), and under the additional assumption (1.33), rates are provided in Theorem 5. Such conditions are verified in the radial case, see Section 9. When it comes to the to convergence of radial derivatives in relative error, i.e. Theorem 6, the results hold up to $p_{Y}$. The second important exponent is $p=\frac{3}{2}$, when the entropy (at least in the form (1.12)) is not defined anymore and we cannot adapt the results of Theorem 5. It is important to notice that $p_{c}$ could be smaller than $\frac{3}{2}$ (in $N=2$), equal to $\frac{3}{2}$ (in $N=3$), or larger than $\frac{3}{2}$ (in $N\in\{4,5\}$), which complicates the graphical representation. We also stress that, for any $D_{2},D_{1}>0$ we have $V_{D_{2}}-V_{D_{1}}\in\mathrm{L}^{1}(\mathbb{R}^{N})$ for any $1<p<2$ and $N\in\{2,3,4,5\}$. In such a case, and under assumption (1.31) the relative entropy ${\mathcal{E}}[v|{V_{D}}]\in\mathrm{L}^{\infty}(\tau_{0},\infty)$ for any $\tau_{0}>0$. The conservation of relative mass, i.e. property (ii) in Theorem 5, is expected to hold under other assumptions of Theorem 5, as it holds for solutions whose initial data is radial-decreasing, see Section 9. It is unknown whether the Fisher information is bounded along the flow, or not. However, this property holds in the case of radially decreasing initial data, see Section 9. Inequality (1.15), i.e. the fact that the Fisher information controls the entropy, can be obtained by comparison with the Hardy–Poincaré inequality, see Proposition 5.9, under the additional assumption (1.33). When $N\geq 6$, the situation is similar but a major difference: the exponents $p_{1}$ and $p_{2}$ appears and $V_{D_{2}}-V_{D_{1}}\in\mathrm{L}^{1}(\mathbb{R}^{N})$ holds only when $p\in\left(1,p_{1}\right)\cup\left(p_{2},2\right)$. Consequently, the conservation of relative mass and the boundedness of the entropy, could hold only in this last range. At the same time, when $p\in\left[p_{1},p_{2}\right]$ the Hardy–Poincaré inequality of Proposition 5.9 does not hold any more. Instead, when $p\in\left(p_{1},p_{2}\right)$ a Hardy-type inequality holds, see [28]. In Table 2 we resume such properties for $N\geq 7$. We notice that the acronym “RDDI” stand for “radial, decreasing whose derivatives satisfy a specific inequality”. Such a class of data will be fully described in Section 9. The case $N=6$ is very similar with the only caveat that $p_{2}=p_{Y}=\frac{3}{2}$.
@@
Range $V_{D_{2}}-V_{D_{1}}\in\mathrm{L}^{1}(\mathbb{R}^{N})$ ? ${\mathcal{E}}[v(\tau)|{V_{D}}]$ ${\mathcal{I}}[v(v)|{V_{D}}]$ $\tfrac{\,{\rm d}}{\,{\rm d}\tau}{\mathcal{E}}[v(\tau)]=-{\mathcal{I}}[v(v)]$ Inequality (1.15) ${\mathcal{I}}[v(v)|{V_{D}}]\geq c\,{\mathcal{E}}[v(\tau)|{V_{D}}]$ Convergence in relative error for radial derivatives
@@ 2 Results in the original variables
[…]
@@ 3.2 Existence and uniqueness
We say that $u$ is a non-negative weak solution to problem (CPLE) on $(0,\infty)\times\mathbb{R}^{N}$ for $1<p<2$ with non-negative initial data ${{u_{0}}}\in\mathrm{L}^{1}_{{\rm loc}}(\mathbb{R}^{N})$ if $u\in\mathrm{L}^{p}((0,\infty);W^{1,p}_{\rm loc}(\mathbb{R}^{N}))\cap C((0,\infty);\mathrm{L}^{1}_{{\rm loc}}(\mathbb{R}^{N}))$ and
@@ 3.3 Boundedness and regularity of solutions
Here we shall mainly focus on the concept of local weak solution (not necessarily non-negative ones) which differ from our definition above by assumptions on the integrability properties of $u$, namely for a solution defined on $(0,T)\times\Omega$ one asks typically that $u\in\mathrm{L}^{\infty}_{\rm loc}((0,T);\mathrm{L}^{2}_{\rm loc}(\Omega))$ and $|\nabla u|\in\mathrm{L}^{p}_{\rm loc}((0,T);\mathrm{L}^{p}_{\rm loc}(\Omega))$. To the best of our knowledge, these kind of solutions have been studied first in [38] where continuity of $\nabla u$ has been proven, with an explicit modulus of continuity, for $p>\max\{1,p_{Y}\}=\max\{1,\frac{2N}{N+2}\}$. In the same range of $p$, the Hölder continuity of the gradient has been obtained in [40] (with some mistakes in the computations, as it was pointed out in [80], which have been fully solved in [39]), see also the review [77]. The threshold value $p_{Y}=\frac{2N}{N+1}$ is sharp for the gradient regularity under no extra assumptions on the solution. Indeed, when $p\leq p_{Y}$ weak solutions are not bounded, see the discussion [42, Section 21.3]. However, in the whole range $1<p<2$ it is known, that bounded weak solutions are Hölder continuous, see [79, 78]. In our case, it is known that solutions to (CPLE) are bounded provided the initial datum ${{u_{0}}}\in\mathrm{L}^{q}_{\rm loc}(\mathbb{R}^{N})$ for $q>N\frac{(2-p)}{p}$, see for instance [43, Section III] and [21, Theorem 2.1]. In conclusion, when $p\in(p_{Y},2)$ and the initial datum is integrable enough, weak solutions to (CPLE) are bounded and, therefore, the function $(t,x)\mapsto u(t,x)$ is $C^{1,\alpha}_{\rm loc}((0,\infty)\times\mathbb{R}^{N})$. However, when $p<p_{c}$, the coefficient $\alpha$ may depend on the function itself other than on $p$ and $N$.
In the present paper, we shall also consider the case $p\leq p_{Y}$, so let us comment on how to obtain the $C^{1,\alpha}_{\rm loc}(\mathbb{R}^{N})$ regularity for solutions in this case. Let us stress, however, that the following reasoning holds in the whole range $1<p<2$. The main idea is to use the concept of viscosity solutions (see [54] for a precise definition). In our case, when a weak solution is continuous, then it is also a viscosity solution, see for instance [54, 71, 46, 45]. Since we shall consider only bounded weak solutions (which are Hölder continuous), the above discussion proves that solutions to problem (CPLE) are, indeed, viscosity solutions. The main advantage of employing this notion of solutions is that, in the last decade, there has been a growing interest in obtaining regularity results for viscosity solutions to equations related to (CPLE), see for instance [46] for a detailed bibliography. For what concerns our investigation, the needed result is [52, Theorem 1.1], where the authors prove the function $(t,x)\mapsto u(t,x)$ is $C^{1,\alpha}_{\rm loc}((0,\infty)\times\mathbb{R}^{N})$ for some $\alpha>0$, which depends on $p$, NN and the solution itself.
@@
Since the uniform convergence in relative error is related to Harnack inequalities, let us conclude this subsection with some considerations on them. The problem of obtaining a precise form of those inequalities has been a long-standing quest. Indeed, in this nonlinear setting, the intrinsic cylinders depend on the solution itself, showing several differences between the case $p_{c}<p<2$ and $1<p\leq p_{c}$. In the good range, we refer to the paper [41], while in the whole range $1<p<2$, it has been proven in [21]. The Harnack inequalities considered in those two papers are valid for local solutions, i.e., no assumption on boundary data is made. For boundary Harnack inequalities, we refer to [62]. Nowadays, several related results are available, see the monograph [41] and references therein.